OpenStudy (abbycross167):

can someone check my work? Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. y = (1/3)x – 1 Steps~ y=-1/3x+1 Let x=0 and solve for y y=-1/3(0)+1 y=0+1 y=1 Plot (0,1) Let x=3 and solve for y y=-1/3(3)+1 y=-1+1 y=0 Plot (3,0) Connect the points and you have

OpenStudy (baru):

y=-1/3x+1 Let x=0 and solve for y y=-1/3(0)+1 y=0+1 y=1 Plot (0,1)

OpenStudy (baru):

That much is correct

OpenStudy (baru):

The rest is wrong

OpenStudy (baru):

It's supposed to be Let y=0 and solve for x

OpenStudy (baru):

Also just check this: question says equation is y=1/3x-1 But ur solution is using y= -1/3x + 1

OpenStudy (abbycross167):

Oh... @baru could you show me how to do this one? please?

OpenStudy (anonymous):

I gave you a medal for helping @baru

OpenStudy (baru):

what is the question? is it y=1/3x -1 or is it y=-1/3x +1 ?

OpenStudy (baru):

thanks @Music2MySoul8 :)

OpenStudy (baru):

@abbycross167

OpenStudy (abbycross167):

y=(1/3)x -1 sorry for not replying sooner... my WiFi is really slow today

OpenStudy (baru):

y=1/3x-1 Let x=0 and solve for y y=1/3(0)-1 y=0-1 y=-1 Plot (0,-1) Let y=0 and solve for x 0=1/3x-1 add 1 to both sides 0+1=1/3x-1+1 1=1/3x multiply both sides by 3 3=3/3x 3=x plot (3,0) connect pts

OpenStudy (abbycross167):

@baru I would put 3 and -1 on my graph?

OpenStudy (baru):

you would put (0,-1) and (3,0)

OpenStudy (baru):



OpenStudy (abbycross167):

Okay thank you soo much you're such a great help!

OpenStudy (baru):

sure :)